if the third of the three consecutive positive odd integers is added to product of the first and second,the result is 144 find the integers
Answer by gc2000(22)

let consecutive positive odd integers = x, x+2, x+4
add 3rd to product of 1st & 2nd, it totals 144:
(x)(x+2) + (x+4) = 144
x^2 + 2x + x + 4 = 144
x^2 + 3x + 4 = 144
x^2 + 3x - 140 = 0
(x + 14)(x - 10) [oops, that doesn't factor, did I make a mistake?]
Check by Trial and error:
9 11 13 ... 9*11 + 13 = 112
11 13 15 ... 11*13 + 15 = 158
No mistake. There is NO SOLUTION [or perhaps the question wasn't worded correctly?]